Prove that the commutative property holds on the matrix multiplication.


1
Expert's answer
2021-10-12T10:19:01-0400

Taking into account that


"\begin{pmatrix}
1 & -1\\
0 & 1
\end{pmatrix}
\begin{pmatrix}
-1 & 1\\
0 & 1
\end{pmatrix}
=
\begin{pmatrix}
-1 & 0\\
0 & 1
\end{pmatrix}," but


"\begin{pmatrix}
-1 & 1\\
0 & 1
\end{pmatrix}
\begin{pmatrix}
1 & -1\\
0 & 1
\end{pmatrix}
=
\begin{pmatrix}
-1 & 2\\
0 & 1
\end{pmatrix},"


we conclude that the commutative property does not hold on the matrix multiplication.